OpenStudy (anonymous):

You have a torn tendon and are facing arthroscopic surgery to fix it. The surgeon explains the risks of the surgery. Infection occurs in 2% of all cases and the repair fails in 11% of the cases. 0.5% of the time the repair fails and infection occurs. What is the probability that the operation is successful and infection-free?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

P(Operation Successful and infection free) = P( (Operation does NOT fail) and (Infection does NOT occur) ) P(Operation Successful and infection free) = P(Operation does NOT fail) * P(Infection does NOT occur) P(Operation Successful and infection free) = (1 - P(Operation fails)) * (1 - P(Infection)) P(Operation Successful and infection free) = (1 - 0.11) * (1 - 0.02) P(Operation Successful and infection free) = (0.89) * (0.98) P(Operation Successful and infection free) = 0.8722 This sounds right, but I'm not sure how/where the 0.5% fits in

OpenStudy (anonymous):

Ok so the answer would be the .8722 part?

jimthompson5910 (jim_thompson5910):

let me draw a probability tree real quick to make sure

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

hmm i can only get part of the tree done, and not the part I want

jimthompson5910 (jim_thompson5910):

maybe a tree won't work here

OpenStudy (anonymous):

Do u think the .87 etc is the right answer

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

yeah i don't have any other reason to doubt it, just not sure how the 0.5% fits in or why they put it there